OpenStudy (anonymous):

hello, could someone please check my statistics work? also, could you tell me the purpose of sample variance and standard deviation? as in, what they show or apply to in the data. thank you.

OpenStudy (amistre64):

variance is the square of the deviation

OpenStudy (amistre64):

deviation tells us how wild a data set is, it is either calm and closely packed, or widely spread out and all over the place

OpenStudy (anonymous):

\[S^2 = (22.5^2 - (22.5^2/30))/30-1\]

OpenStudy (anonymous):

16.875?

OpenStudy (amistre64):

i assume theres a data set you are working with?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

can you post it? or a screenshot/pic?

OpenStudy (anonymous):

yeah, hold on

OpenStudy (amistre64):

\[var(x)=\frac{\sum(x-\bar x)^2}{N-1}\] \[s=\sqrt{var}\]

OpenStudy (anonymous):

how do you extract an observation from that result, though? what am i looking for in terms of variance or deviation, respectively?

OpenStudy (amistre64):

i see you did a stem and leaf, and the data appears to be of a normal distribution. high in the middle and small at the ends, and symmetric

OpenStudy (anonymous):

so what does that say about the deviation?

OpenStudy (anonymous):

actually i just noticed in my notes that my professor says the standard deviation is the square root of the variance...ack what am i mixed up with

OpenStudy (amistre64):

variance is a self serving number, it is the sum of the squared distances from the mean.